In a normal distribution with µ = $19,000 and the Interquartile range is $5000 (IQR=5000), find Q1.
Answer #1

Answer)

In normal distribution

Q1 = mean - 0.67*s.d..... EQUATION 1

Q3 = mean + 0.67*s.d.... EQUATION 2

Iqr = Q3 - Q1 = 5000

Subtracting equation 1 from 2

We get,

Q3 - Q1 = 1.34*s.d

5000 = 1.34*s.d {as Q3 - Q1 = 5000}

S.d = 3731.34328358

So, Q1 = 19000 - 0.67*3731.34328358 = 16500

Answer is 16500